# Measuring the Measures: Discriminative Capacity of Representational Similarity Metrics Across Model Families

### v4

@@ 2.1 Model Selection and Dataset
We analyze 35 models across four primary categories: supervised CNNs, self-supervised CNNs, supervised Transformers, and self-supervised Transformers. We treat ConvNeXt convnet and Swin swin as distinct families due to their hybrid nature—ConvNeXt incorporates Transformer-inspired design principles within a convolutional architecture, while Swin introduces CNN-like inductive biases into the Transformer framework. For evaluation, we use a curated subset of ImageNet-1k imagenet_cvpr09 . Complete model specifications and dataset details are provided in Appendices A and B.
@@ 2.2 Metrics for Representational Similarity
Soft Matching (SoftMatch) softmatching .
SoftMatch finds an optimal probabilistic correspondence between units in two representations by solving a transport problem. The method seeks a mapping matrix $\mathbf{M}\in\mathcal{T}(N_{i},N_{j})$ in the transportation polytope, where each entry $M_{ij}$ represents the matching weight between units. The constraints ensure doubly-stochastic normalization: $\sum_{j=1}^{N_{j}}M_{ij}=\frac{1}{N_{i}},\quad\sum_{i=1}^{N_{i}}M_{ij}=\frac{1}{N_{j}}$ The optimal mapping minimizes the reconstruction error $\|\mathbf{X}_{j}-\mathbf{M}\mathbf{X}_{i}\|_{2}^{2}$ and is solved using the network simplex algorithm. When $N_{i}=N_{j}$, this reduces to finding an optimal permutation. The final alignment score is the correlation between $\mathbf{X}_{j}$ and the transformed representation $\mathbf{M}\mathbf{X}_{i}$.
@@ 3 Results
Our systematic evaluation reveals clear differences in the discriminative capacity of representational similarity metrics (Figures 1, 2). RSA demonstrates the strongest overall separability with a d-prime of 3.95, closely followed by SoftMatch ($d^{\prime}=3.59$), then Procrustes ($d^{\prime}=3.03$), with Linear Predictivity showing substantially weaker discrimination ($d^{\prime}=2.02$). The pattern is particularly pronounced for silhouette coefficients, where RSA achieves a notably high score of 0.56—indicating robust within-family clustering—while the mapping-based metrics show progressively weaker clustering: SoftMatch (0.29), Procrustes (0.21), and Linear Predictivity (0.13). ROC-AUC analysis confirms this hierarchy (Figure 3: RSA (0.9257) and SoftMatch (0.8994) achieve superior classification of within- versus between-family pairs, followed by Procrustes (0.8979) and Linear Predictivity (0.8137).
@@ Appendix A Experiment Settings
We use ImageNet-1K [9]. To control class imbalance and reduce compute, we evaluate on a balanced subset drawn from the test/validation split: 50 images per class sampled uniformly at random (fixed seed). All representational metrics are computed on this subset.

### v5

@@ 2.1 Model Selection and Dataset
We analyze 35 models across four primary categories: supervised CNNs, self-supervised CNNs, supervised Transformers, and self-supervised Transformers. We treat ConvNeXt convnet and Swin swin as distinct families due to their hybrid nature—ConvNeXt incorporates Transformer-inspired design principles within a convolutional architecture, while Swin introduces CNN-like inductive biases into the Transformer framework. For evaluation, we use the test set of ImageNet-1k imagenet_cvpr09 . Complete model specifications and dataset details are provided in Appendices A and B.
@@ 2.2 Metrics for Representational Similarity
Soft Matching (softmatching, ).
Soft Matching (SoftMatch) generalizes permutation distance (ding2021grounding, ) to representations with different numbers of units by relaxing permutations to “soft permutations.” Specifically, The method seeks a mapping matrix $\mathbf{P}\in\mathcal{T}(N_{i},N_{j})$ in the transportation polytope (de2013combinatorics, ), where each entry $P_{ij}$ represents the matching weight between units. The constraints ensure doubly-stochastic normalization: $\sum_{j=1}^{N_{j}}P_{ij}=\frac{1}{N_{i}},\quad\sum_{i=1}^{N_{i}}P_{ij}=\frac{1}{N_{j}}$. The optimization problem is
$d_{T}(\mathbf{X}_{i},\mathbf{X}_{j})=\min_{\mathbf{P}\in\mathcal{T}(N_{i},N_{j})}\sum_{k,l}\mathbf{P}_{kl}\,\lVert x_{i}^{(k)}-x_{j}^{(l)}\rVert^{2},$
where $x_{i}^{(k)}$ and $x_{j}^{(l)}$ are the $k$-th and $l$-th columns (units) of $\mathbf{X}_{i}$ and $\mathbf{X}_{j}$. The optimal transport plan $\mathbf{P}^{\star}$ is found via the network simplex algorithm. When $N_{i}=N_{j}$, this reduces to an optimal permutation. The final similarity score is the mean unit-wise correlation between $\mathbf{X}_{j}$ and $\mathbf{X}_{i}\mathbf{P}^{\star}$.
@@ 3 Results
Our systematic evaluation reveals clear differences in the discriminative capacity of representational similarity metrics (Figures 1, 2). RSA demonstrates the strongest overall separability with a d-prime of 3.95, closely followed by SoftMatch ($d^{\prime}=3.59$), then Procrustes ($d^{\prime}=2.96$), with Linear Predictivity showing substantially weaker discrimination ($d^{\prime}=2.09$). The pattern is particularly pronounced for silhouette coefficients, where RSA achieves a notably high score of 0.56—indicating robust within-family clustering—while the mapping-based metrics show progressively weaker clustering: SoftMatch (0.30), Procrustes (0.22), and Linear Predictivity (0.14). ROC-AUC analysis confirms this hierarchy (Figure 3: RSA (0.9257) achieves superior classification of within- versus between-family pairs, while SoftMatch (0.8989) and Procrustes (0.9004) obtain very similar ROC-AUCs, both outperforming Linear Predictivity (0.8273).
@@ Appendix A Experiment Settings
To control class imbalance and reduce compute, we use the ImageNet-1K [9] validation set, which is balanced with 50 images per class. All representational metrics are computed on this dataset.